OpenStudy (brittanyrolfe):

A 25.9 g sample of solid CO2 (dry ice) is added to a container at a temperature of 100 K with a volume of 4.8 L . If the container is evacuated (all of the gas removed), sealed and then allowed to warm to room temperature T = 298 K so that all of the solid CO2 is converted to a gas, what is the pressure inside the container? Express your answer to two significant figures and include the appropriate units.

OpenStudy (brittanyrolfe):

Use the equation: \[\frac{ P _{1} V _{1} }{ n _{1}T _{1} }=\frac{ P _{2}V _{2} }{ n _{2}T _{2} }\]

OpenStudy (brittanyrolfe):

@Jamierox4ev3r

OpenStudy (jiteshmeghwal9):

As the container is evacuated so at the room temperature the whole pressure will be due to 25.9 gm CO2 gas which is equivalent to \(\frac{25.9}{44}\) moles of CO2. the volume occupied by the gas will be the volume f the container hich is 4.8L & temperature is 298 K. Plg these values in gas equation \[PV=nRT\] where P is pressure V is volume occupied n is numbr of moles of CO2 R = Gas constant = 0.082 T=298 K *u will gt pressure in atm.

OpenStudy (caozeyuan):

Maybe I am thinking too deep, but lets assume the solid CO2 is in equilibrium with its own gas vapor due to evaporation \[CO2_{(s)}\leftarrow \rightarrow CO2_{(g)}\]

OpenStudy (caozeyuan):

now if we vacuum the container, CO2(g) will be removed, but CO2(s) will not, therefore, by Le Chatlier Principle, CO2(s) will convert to CO2(g) and get sucked into the vacuum.

OpenStudy (caozeyuan):

So does this happen or not, why?

OpenStudy (brittanyrolfe):

Thank you @jiteshmeghwal9 I thought, since it gave the before conditions that we were to use the equation I had posted. But the steps you provided worked. :)